OpenStudy (unikornkiller):

A salesperson works 40 hours per week at a job where she has two options for being paid. Option A is an hourly wage of $24. Option B is a commission rate of 10% on weekly sales. How much does she need to sell this week to earn the same amount with the two options?

OpenStudy (unikornkiller):

help

OpenStudy (gorv):

let she sales =x

OpenStudy (gorv):

on hourly basis income=40*24=960

OpenStudy (gorv):

by sale =x*10/100=x/10

OpenStudy (gorv):

both should be equal

OpenStudy (gorv):

960=x/10 multiply both side by 10 thtazz it